# On Hardness and Approximation of Broadcasting in Sparse Graphs

### v1

@@ 4.1.1 PTAS for Prefix Covering
Given two sets $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering set $\mathfrak{R}_{p}(C)$ (if one exists) or correctly reports that no such covering is possible.
Proof. Both the ground set $\mathfrak{R}_{p}(S)$ and the covering set $\mathfrak{R}_{p}(C)$ have constant support $p$. Fix a distinct value $v\in\mathfrak{R}_{p}(S)$. For each pair $(a,b)$ of elements from $\mathfrak{R}_{p}(C)$, we can encode how many times $(a,b)$ is used to cover elements of $\mathfrak{R}_{p}(S)$ equal to $v$. Since there are at most $|S|$ such elements, this count can be represented in $\log|S|$ bits. As there are at most $p^{2}$ pairs from $\mathfrak{R}_{p}(C)$ and $p$ distinct values in $\mathfrak{R}_{p}(S)$, any candidate covering can be encoded using at most $p^{3}\log|S|$ bits. Hence, the number of candidate solutions is bounded by $\mathcal{O}(2^{p^{3}}|S|)$. Checking whether a candidate encoding yields a valid covering takes $\mathcal{O}(|S|+|C|)$ time. Therefore, the total running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+|C|))$. Moreover, whenever a valid covering exists, the algorithm outputs it by interpreting the corresponding encoding; otherwise, it correctly reports that no covering is possible. $\Box$
@@
To summarize, given a covering guess size $m$ and ground set $\mathfrak{R}_{p}(S)$, we form $\mathfrak{R}_{p}(C)$, and check whether we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ with an exhaustive approach (Lemma 4.5). If we cannot find a covering, we return bad-guess; we cannot cover $\mathfrak{R}_{p}(S)$ using $[m]$. If we find a solution, use the covering given by $\mathfrak{R}_{p}(C)$ to cover $\mathfrak{R}_{p}(S)$ using $m^{\prime}]$, where $m^{\prime}=m+\lceil m/p\rceil$. This gives a $(1+1/p)$-approximation for covering $\mathfrak{R}_{p}(S)$. Applying the Proposition 4.4, we can get an approximating factor of $(1+2/p)(1+1/p)$, which can be as close as we want to $1$ by setting $p$ large enough. Using binary search adds a multiplicative factor of $\log\boldsymbol{OPT}$ to the runtime we used for the exhaustive search (see Lemma 4.5). Hence, the overall running time is $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$. We can conclude the following:
@@
There is a PTAS for the Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
@@
    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.1. Given that $C^{\prime}=[m^{\prime}]$ covers $\mathfrak{R}_{p}(S)$, it covers $S$ as well. Using Proposition 4.4, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)(1+2/p)\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
Lemma 4.2 and Theorem 4.6 imply the following Corollary.
@@
There exists a PTAS for Telephone Broadcasting in $k$-cycle graphs. In particular, for a $k$-cycle graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
4.2 PTAS for $k$-Path Graphs
In this section, we present a polynomial-time approximation scheme (PTAS) for Telephone Broadcasting in $k$-path graphs. Consider a $k$-path graph $G$ with endpoints $s$ and $t$.
@@ 4.2 PTAS for $k$-Path Graphs
In what follows, we assume that $s$ is the designated source, holding the message at time 0, and that $t$ acts as a late source, receiving the message at time $\alpha$, with $\alpha\leq d(s,t)$, where $d(s,t)$ denotes the distance between $s$ and $t$ in $G$. We show that any instance of the (single-source) problem with an arbitrary originator can be reduced to this double-source setting (with $s$ as the source and $t$ as the late source) without changing the approximation factor (Lemma 4.8).
@@
In Lemma 4.10, we establish an equivalence between double-source broadcasting and the integer covering problem that we call Double Prefix Covering (Definition 4.9).
@@
Finally, by applying a rounding-and-covering approach similar to the one used for Prefix Covering in $k$-cycle graphs, we obtain a PTAS for Double Prefix Covering, and consequently for Telephone Broadcasting in $k$-path graphs (Section 4.2.1). In what follows, we describe these steps in detail. First, we prove the following lemma, which implies we can restrict our attention to the double-source broadcasting.
@@
Any $c$-approximation algorithm for Telephone Broadcasting in $k$-path graphs in the double-source setting—where $s$ is the primary source and $t$ is a late source receiving the message at time $\alpha$—can be extended to a $c$-approximation algorithm for the (single-source) general setting, where the originator $v_{o}$ may be any vertex of the graph.
@@
Proof. Let $v_{o}$ be the originator vertex lying in path $P_{i}$ at distances $d(v_{o},s)$ and $d(v_{o},t)$ from endpoints $s$ and $t$, respectively. Without loss of generality, assume $d(v_{o},s)\leq d(v_{o},t)$. Let $d(v_{o},s)$. The message is first forwarded along the shortest path from $v_{o}$ to $s$, which takes $d(v_{o},s)$ rounds. At round $d(v_{o},s)+1$, vertex $s$ is informed and can begin broadcasting. Vertex $t$ is informed either directly through $v_{o}$ (after $d(v_{o},t)$ rounds) or through $s$ via some other path $P_{j}$ (whichever occurs first). Let $\alpha$ denote the delay with which $t$ receives the message relative to $s$. Once both $s$ and $t$ are informed, they proceed with the $c$-approximation algorithm on the remaining graph.
@@
We introduce the following problem, which is the key part of our PTAS algorithms.
@@
The input to the Double Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers together with a non-negative integer $\beta$. The objective is to determine the minimum integer $m$ such that there exist functions $c:S\to\{0\}\cup[m]$ and $d:S\to\{0\}\cup[m-\beta]$ satisfying $c(i)+d(i)\geq\ell_{i}$ for all $i\in[n]$. Moreover, injectivity is required in the following sense: for any $i\neq j$, if both $c(i)>0$ (respectively $d(i)>0$), then $c(i)\neq c(j)$ (respectively $d(i)\neq d(j)$). In this setting, we say that each $\ell_{i}$ is covered by the pair $\{c(i),d(i)\}$, and we refer to $C=[m]$ and $D=[m-\beta]$ as the covering sets of $S$.
@@
When $S=\{8,4,4,2\},$ the optimal solution is given by $C=[5]$ and $\beta=2$, and we have $c(1)=5,d(1)=3$; $c(2)=4,d(2)=0$, $c(3)=2,d(3)=2$, $c(4)=1,d(4)=1$.
In what follows, we establish the equivalence between double-source broadcasting in $k$-path graphs and Double Prefix Covering (Definition 4.9).
@@
Let $G$ be a $k$-path graph formed by $k$ paths with common endpoints $s$ and $t$, where the $i$th path contains $\ell_{i}\geq 0$ internal vertices (in addition to $s$ and $t$). Consider a double-source broadcasting instance on $G$, where $s$ receives the message at time 0 and $t$ at time $\alpha$. Now, define an instance $I$ of Double Prefix Covering as $I=(S=\{\ell_{1},\ldots,\ell_{k}\},\;\beta=\alpha-1)$. Then computing the minimum broadcast time for this double-source Telephone Broadcasting instance is equivalent to solving the Double Prefix Covering instance $I$.
@@
4.2.1 PTAS for Double Prefix Covering
In the light of Lemmas 4.8 and 4.10, presenting a PTAS for broadcasting in $k$-path graphs reduces to presenting a PTAS for the Double Prefix Covering problem. This is the focus of the present section. Specifically, we prove Theorem 4.13, which establishes the existence of a $(1+\epsilon)$-approximation algorithm for Double Prefix Covering for any $\epsilon>0$. Our approach closely parallels the rounding and covering techniques used in Section 4.1.1 for $k$-cycle graphs. In particular, it also addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the two covering sets into simplified versions with constant support (i.e., bounded numbers of distinct elements), so that exhaustive search becomes computationally feasible while still preserving provable approximation guarantees.
Our PTAS for Double Prefix Covering proceeds by reducing the original instance to one with bounded support. Given an input multiset $S$, we first form its rounded version $\mathfrak{R}_{p}(S)$, where each element is rounded up to a larger element of $S$ so that the number of distinct values becomes $p$. Here $p$ depends only on $\epsilon$. Although rounding inflates item sizes, we show that the optimal covering cost for $\mathfrak{R}_{p}(S)$ is at most a $(1+1/p)$ factor larger than that of $S$ (Proposition 4.11). Thus, solving the rounded instance suffices for approximation.
To search for the optimal covering, we use binary search over candidate values $m$, where $m$ determines the covering sets $C=[m]$ and $D=[m-\beta]$. For each candidate $m$, we construct rounded sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, each with support size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, then the original instance can be covered using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ with $m^{\prime}\leq(1+1/p)m$ (4.2). Conversely, if we cannot cover $\mathfrak{R}_{p}(S)$ with the rounded sets, then it is not possible to cover it using $[m]$ and $m-\beta]$. It therefore suffices to test whether $\mathfrak{R}_{p}(S)$ is coverable using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$. Since all three sets have constant support (bounded by $p$), this check can be carried out by exhaustive search in constant time (Lemma 4.12). The remainder of this section provides details of the algorithm.
@@ 4.2.1 PTAS for Double Prefix Covering
Since $\boldsymbol{OPT}(S)\geq|S|$, it follows that any covering of $\mathfrak{R}_{p}(S)$ yields a $(1+1/p)$-approximation for covering $S$. This allows us to restrict attention to $\mathfrak{R}_{p}(S)$, which has constant support $p$.
To cover $\mathfrak{R}_{p}(S)$, we apply a binary strategy. For a candidate value $m$, we design a procedure that either produces a covering using $m+\lceil m/p\rceil]$, or certifies that no covering exists with $[m]$. By iteratively refining $m$, we identify the smallest value for which $\mathfrak{R}_{p}(S)$ can be covered using the rounded covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, where $C=[m]$ and $D=[m-\beta]$.

### v2

@@ 4.1 EPTAS for Cycle-star Graphs
Given two sets $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, there exists an algorithm that, in time $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ (if one exists) or correctly reports that no such covering is possible.
Proof sketch: We model the covering problem as a feasibility instance of integer linear programming with one variable for each possible way of covering a distinct element in $\mathfrak{R}_{p}(S)$ using one or two elements in $\mathfrak{R}_{p}(C)$. Given that both the ground multiset $\mathfrak{R}_{p}(S)$ and the covering multiset $\mathfrak{R}_{p}(C)$ have constant support $p$, there are only $\mathcal{O}(p^{3})$ such variables. The constraints enforce that (i) every ground element is covered exactly once, (ii) no covering type is used more than its multiplicity, and (iii) only feasible type combinations are allowed. By Lenstra’s algorithm, the resulting integer linear feasibility can be solved in time $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$. Thus, we can either find a feasible covering or correctly conclude that none exists. $\Box$
Proof. Let $V$ and $T$ be the set of distinct elements in $\mathfrak{R}_{p}(S)$ and $\mathfrak{R}_{p}(C)$, respectively. Then $|V|\leq p$ and $|T|\leq p$. For each pair $\{i,j\}$ of items in $T$ and each ground type $v\in V$, introduce an integer variable $x_{i,j,v}\in\mathbb{Z}_{\geq 0}$, intended to denote how many times the pair $(i,j)$ is used to cover elements of type $v$. Similarly, for each element $i\in T$ and ground type $v\in V$, we let $x_{i,v}\in\mathbb{Z}_{\geq 0}$ to denote the number of times $i$ is used to cover elements of type $v$. There are $\big({{|T|}\choose{2}}+|T|\big)|V|\leq p^{3}$ variables.
We write a feasibility ILP whose constraints enforce that (i) every ground element is covered exactly once, and (ii) cover-items are not used more than their multiplicities.
Cover all ground elements. For each $v\in V$, we let $\mu_{S}(v)$ be the multiplicity of $v$, that is, $\mu_{S}(v):=|\{s\in\mathfrak{R}_{p}(S):s=v\}|$. To ensure $v$ is covered exactly once, we add a constraint:
$\sum_{i,j\in T}x_{i,j,v}+\sum_{i\in T}x_{i,v}\;=\;\mu_{S}(v).$ (1)
Do not overuse cover-items. For each $t\in T$, let $\mu_{C}(t)$ be the multiplicity of type $t$ in $\mathfrak{R}_{p}(C)$, that is, $\mu_{C}(t):=|\{s\in\mathfrak{R}_{p}(C):s=t\}|$. Each time we use a pair $(i,j)$ we consume one item of type $i$ and one item of type $j$ (with the understanding that if $i=j$ then we consume two copies). Thus, for each $t\in T$ we impose
$\sum_{v\in V}\ \big(\sum_{i,j\in T}\bigl(\mathbf{1}[i=t]+\mathbf{1}[j=t]\bigr)\,x_{i,j,v}+\sum_{i\in T}\bigl(\mathbf{1}[i=t]\big)\,x_{i,v}\bigr)\;\leq\;\mu_{C}(t).$ (2)
Here, $\mathbf{1}[j=t]$ denotes the indicator function, which is equal to $1$ if $j=t$ and 0 otherwise.
Type-feasibility. Finally, we forbid pairs that cannot cover type $v$ by setting $x_{i,j,v}=0$ whenever $(i,j)$ is not a valid set for covering $v$, that is, $i+j<v$. Similarly, we force $x_{i,v}=0$ when $i<v$.
This ILP has at most $p^{3}$ integer variables, and the number of constraints is $\mathcal{O}(p^{3})$.
By Lenstra’s theorem [22] (and subsequent improvements by Kannan [21]), integer linear feasibility with a fixed number $d$ of variables can be solved in time $d^{\mathcal{O}(d)}\cdot\operatorname{poly}(L)$, where $L$ is the encoding length of the ILP. Here $d\leq p^{3}$ and $L=\operatorname{poly}(|S|+|C|+p)$, hence the feasibility test runs in $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|)$ time. If feasible, the resulting integer solution specifies a valid covering; otherwise, no such solution exists. $\Box$
@@
To summarize, given a covering guess size $m$ and ground set $\mathfrak{R}_{p}(S)$, we form $\mathfrak{R}_{p}(C)$, and check whether we can cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$ with a feasibility ILP (Lemma 4.5). If we cannot find a covering, we return bad-guess; we cannot cover $\mathfrak{R}_{p}(S)$ using $[m]$. If we find a solution, use the covering given by $\mathfrak{R}_{p}(C)$ to cover $\mathfrak{R}_{p}(S)$ using $m^{\prime}]$, where $m^{\prime}=m+\lceil m/p\rceil$. This gives a $(1+1/p)$-approximation for covering $\mathfrak{R}_{p}(S)$. Applying the Proposition 4.4, we can get an approximating factor of $(1+2/p)(1+1/p)$, which can be as close as we want to $1$ by setting $p$ large enough. Using binary search adds a multiplicative factor of $\log\boldsymbol{OPT}$ to the runtime we used for ILP (see Lemma 4.5). Hence, the overall running time is $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$.
We can conclude the following:
@@
There is a $(1+\epsilon)$-approximation for the Prefix Covering problem that terminates in $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
@@
    Use the feasibility ILP (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.1. Given that $C^{\prime}=[m^{\prime}]$ covers $\mathfrak{R}_{p}(S)$, it covers $S$ as well. Using Proposition 4.4, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)(1+2/p)\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$ time, giving total complexity $p^{O(p^{3})}\cdot\operatorname{poly}(|S|+\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
EPTAS for Telephone Broadcasting in Cycle-Star graphs
Given the equivalence between Telephone Broadcasting in cycle-star graphs and the Prefix Covering problem, as established in Lemma 4.2, along with the EPTAS presented in Theorem 4.6 for the Prefix Covering problem, we can conclude the following.
@@
There exists an EPTAS for Telephone Broadcasting in cycle-star graphs. In particular, for a cycle-star graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
Extension to multi-source broadcasting
The EPTAS of Corollary 4.7 extends naturally to the multi-source broadcasting setting. First, we provide a sketch of this extension.
The main idea is to guess the optimal broadcast time $\boldsymbol{br}$ and, for each guess, attempt to construct a broadcast schedule. If $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, the procedure returns a schedule of length at most $(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1$. Trying all relevant guesses for $\boldsymbol{br}$ and selecting the shortest resulting schedule yields an EPTAS.
For each fixed guess $\boldsymbol{br}$, we reduce the multi-source instance to a single-source instance as follows. We first perform a preprocessing step that removes a window of at most $\boldsymbol{br}$ vertices around each non-center source. These vertices can always be informed within $\boldsymbol{br}+1$ rounds directly from their respective sources. If this preprocessing disconnects the graph or leaves a vertex at a distance larger than $\boldsymbol{br}$ from every source, we reject the guess value $\boldsymbol{br}$. Otherwise, the remaining graph $G^{\prime}$ is a graph of distance 1 to a path forest (See Figure 1). This graph has the center as the unique bottleneck through which all remaining vertices must be informed. We therefore reduce the problem to a single-source instance $(G^{\prime},s)$, where $s$ is the center. This instance is transformed into a modified version of the Prefix Covering problem in which cycle components allow one or two-element covers as before, while path components (created by preprocessing) correspond to ground elements that must be covered by a single item. The rounding and approximation steps from the single-source EPTAS remain unchanged, and the resulting rounded instance of Prefix Covering has constant support. It can therefore be solved exactly using a fixed-dimension ILP as before.
Now, we present details on the extension of the EPTAS of Corollary 4.7 to the multi-source broadcasting setting.
Suppose we are given an instance $(G,S)$, where $S$ is the set of all sources, and a fixed parameter $\epsilon>0$. Let $\boldsymbol{br}$ be a broadcast-time guess value. We describe a procedure that, when $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, outputs a valid broadcasting scheme that completes within $(1+\epsilon)\boldsymbol{br}+1$. Otherwise, it returns either Bad-Guess or a scheme with no guarantee on its broadcast time. As we do not know the optimal broadcast time, we iterate over all $\boldsymbol{br}\in[\lceil\log n\rceil,n]$ and take the scheme with the best broadcasting time among all returned schemes. This way, we can achieve broadcast time $(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1$.
Preprocessing. For each non-center source $s_{i}$, we remove all vertices that lie within distance $\boldsymbol{br}$ of $s_{i}$ on either of the two paths connecting $s_{i}$ to the center, excluding the center itself. (The vertex $s_{i}$ is removed as well.) If the center is not initially a source, let $s_{m}$ be the source closest to the center, at distance $d_{\min}$, and remove up to $\boldsymbol{br}-d_{\min}$ additional vertices along the remaining path between the center and $s_{m}$.
Let $G^{\prime}$ denote the remaining graph. If $G^{\prime}$ is disconnected, or if $d_{\min}>\boldsymbol{br}$, then there exists a vertex whose distance from every source exceeds $\boldsymbol{br}$, and we correctly return Bad-Guess. Otherwise, $G^{\prime}$ will be a graph of distance $1$ to path forests (see Figure 1).
The preprocessing removes a window of at most $\boldsymbol{br}$ vertices around each non-center source. Note that all removed vertices can be informed within $\boldsymbol{br}+1$ rounds by broadcasting from their respective sources.
Reduction to covering. Consider an instance $(G^{\prime},s)$ of single-source Telephone Broadcasting with the center being the source. We associate $(G^{\prime},s)$ with an instance $\mathcal{S}$ of the Prefix Covering problem, defined analogously to Lemma 4.2 for cycle-star graphs. The only difference is that some elements of the ground multiset $\mathcal{S}$, which we refer to as path elements, correspond to paths in $G^{\prime}$ rather than cycles. Each path element must be covered by a subset of size exactly $1$ (it cannot be covered with two elements from the ground set).
This modified covering instance can be solved in the same way as discussed in Section 4.1. In particular, the rounding steps remain unchanged. The only additional constraint arises in the final step (Lemma 4.5), where we ensure that each path element in the rounded ground set is covered by a single-item subset. This produces a broadcast schedule for $(G^{\prime},s)$ that completes within $(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)$ rounds.
Combining the solutions. Overall, broadcasting in the original multi-source instance $G$ completes within
$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}.$
The term $\boldsymbol{br}+1$ accounts for informing all vertices removed during the preprocessing step. The second term captures the time required to inform the center and its neighbor along the path toward $s_{m}$ (namely, $d_{\min}+1$ rounds), followed by the execution of the single-source broadcast within $G^{\prime}$.
Suppose $br={\boldsymbol{br}^{*}}(G,S)$. Now, if $G^{\prime}$ has no vertices other than the center, then we have found a broadcasting schedule with $\boldsymbol{br}+1={\boldsymbol{br}^{*}}(G,S)+1$ rounds. Otherwise, in the optimal scheme, vertices in $G^{\prime}$ must be informed through the center. This is because the vertices in $G^{\prime}$ are the ones that were not removed when forming $G^{\prime}$ and thus are at a distance further than $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$ from the sources in their respective cycles. Therefore, as the center cannot be informed earlier than round $d_{\min}$, we can write ${\boldsymbol{br}^{*}}(G,S)\geq d_{\min}+{\boldsymbol{br}^{*}}(G^{\prime},s)$. Hence, the broadcast time of our scheme is at most
$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}\leq(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1.$
4.2 EPTAS for Melon Graphs
@@
In this section, we present an efficient polynomial-time approximation scheme (EPTAS) for Telephone Broadcasting in melon graphs. Consider a melon graph $G$ with endpoints $s$ and $t$. In what follows, we assume that $s$ is the designated source, holding the message at time 0, and that $t$ acts as a late source, receiving the message at time $\alpha$, with $\alpha\leq d(s,t)$, where $d(s,t)$ denotes the distance between $s$ and $t$ in $G$. We show that any instance of the (single-source) problem with an arbitrary originator can be reduced to this double-source setting (with $s$ as the source and $t$ as the late source) without changing the approximation factor (Lemma 4.8).
@@
In Lemma 4.10, we establish an equivalence between double-source Telephone Broadcasting and the integer covering problem that we call Double Prefix Covering (Definition 4.9).
@@
Finally, by applying a rounding-and-covering approach similar to the one used for Prefix Covering in cycle-star graphs, we obtain an EPTAS for Double Prefix Covering, and consequently for Telephone Broadcasting in melon graphs (Section 4.2). In what follows, we describe these steps in detail. First, we prove the following lemma, which implies we can restrict our attention to the double-source Telephone Broadcasting.
@@
Any $c$-approximation algorithm for Telephone Broadcasting in melon graphs in the double-source setting—where $s$ is the primary source and $t$ is a late source receiving the message at time $\alpha$—can be extended to a $c$-approximation algorithm for the (single-source) general setting, where the originator $v_{o}$ may be any vertex of the graph.
@@
Proof. Let $v_{o}$ be the originator vertex lying in path $P_{i}$ at distances $d(v_{o},s)$ and $d(v_{o},t)$ from endpoints $s$ and $t$, respectively. Without loss of generality, assume $d(v_{o},s)\leq d(v_{o},t)$. The message is first forwarded along the shortest path from $v_{o}$ to $s$, which takes $d(v_{o},s)$ rounds. At round $d(v_{o},s)+1$, vertex $s$ is informed and can begin broadcasting. Vertex $t$ is informed either directly through $v_{o}$ (after $d(v_{o},t)$ rounds) or through $s$ via some other path $P_{j}$ (whichever occurs first). Let $\alpha$ denote the delay with which $t$ receives the message relative to $s$. Once both $s$ and $t$ are informed, they proceed with the $c$-approximation algorithm on the remaining graph.
@@
From Telephone Broadcasting in Melon Graphs to Double Prefix Covering
We introduce the following problem, which is the key part of our EPTAS.
@@
The input to the Double Prefix Covering problem is a ground multiset $S=\{\ell_{1},\ell_{2},\ldots,\ell_{n}\}$ of positive integers together with a non-negative integer $\beta$. The objective is to determine the minimum integer $m$ such that there exist functions $c:S\to\{0\}\cup[m]$ and $d:S\to\{0\}\cup[m-\beta]$ satisfying $c(i)+d(i)\geq\ell_{i}$ for all $i\in[n]$. Moreover, injectivity is required in the following sense: for any $i\neq j$, if both $c(i)>0$ (respectively, $d(i)>0$), then $c(i)\neq c(j)$ (respectively $d(i)\neq d(j)$). In this setting, we say that each $\ell_{i}$ is covered by the pair $\{c(i),d(i)\}$, and we refer to $C=[m]$ and $D=[m-\beta]$ as the covering sets of $S$.
@@
When $S=\{8,4,4,2\}$ and $\beta=2$, the optimal solution is given by $C=[5]$, and we have $c(1)=5,d(1)=3$; $c(2)=4,d(2)=0$, $c(3)=2,d(3)=2$, $c(4)=1,d(4)=1$.
In what follows, we establish the equivalence between double-source Telephone Broadcasting in melon graphs and Double Prefix Covering (Definition 4.9).
@@
Let $G$ be a melon graph formed by $k$ paths with common endpoints $s$ and $t$, where the $i$th path contains $\ell_{i}\geq 0$ internal vertices (in addition to $s$ and $t$). Consider a double-source Telephone Broadcasting instance on $G$, where $s$ receives the message at time 0 and $t$ at time $\alpha$. Now, define an instance $I$ of Double Prefix Covering as $I=(S=\{\ell_{1},\ldots,\ell_{k}\},\;\beta=\alpha-1)$. Then computing the minimum broadcast time for this double-source Telephone Broadcasting instance is equivalent to solving the Double Prefix Covering instance $I$.
@@
EPTAS for Double Prefix Covering
In the light of Lemmas 4.8 and 4.10, presenting an EPTAS for Telephone Broadcasting in melon graphs reduces to presenting an EPTAS for the Double Prefix Covering problem. This is the focus of the present section. Specifically, we prove Theorem 4.13, which establishes the existence of a $(1+\epsilon)$-approximation algorithm for Double Prefix Covering for any $\epsilon>0$. Our approach closely parallels the rounding and covering techniques used in Section 4.1 for cycle-star graphs. In particular, it also addresses the covering problem through a two-stage approximation scheme that leverages the concept of rounded sets (see Definition 4.3) to reduce the problem complexity. The main idea is to transform the original ground multiset and the two covering sets into simplified versions with constant support (i.e., bounded numbers of distinct elements), so that an ILP formulation becomes computationally feasible while still preserving provable approximation guarantees.
Our EPTAS for Double Prefix Covering proceeds by reducing the original instance to one with bounded support. Given an input multiset $S$, we first form its rounded version $\mathfrak{R}_{p}(S)$, where each element is rounded up to a larger element of $S$ so that the number of distinct values becomes $p$. Here $p$ depends only on $\epsilon$. Although rounding inflates item sizes, we show that the optimal covering cost for $\mathfrak{R}_{p}(S)$ is at most a $(1+1/p)$ factor larger than that of $S$ (Proposition 4.11). Thus, solving the rounded instance suffices for approximation.
To search for the optimal covering, we use binary search over candidate values $m$, where $m$ determines the covering sets $C=[m]$ and $D=[m-\beta]$. For each candidate $m$, we construct rounded sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, each with support size $p$. We prove that if $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, then the original instance can be covered using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ with $m^{\prime}\leq(1+1/p)m$ (4.2). Conversely, if we cannot cover $\mathfrak{R}_{p}(S)$ with the rounded sets, then it is not possible to cover it using $[m]$ and $m-\beta]$. It therefore suffices to test whether $\mathfrak{R}_{p}(S)$ is coverable using $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$. Since all three sets have constant support (bounded by $p$), this check can be carried out by an ILP formulation in polynomial time (Lemma 4.12). The remainder of this section provides details of the algorithm.
@@
Since $\boldsymbol{OPT}(S)\geq|S|$, it follows that any covering of $\mathfrak{R}_{p}(S)$ yields a $(1+1/p)$-approximation for covering $S$. This allows us to restrict attention to $\mathfrak{R}_{p}(S)$, which has constant support $p$. To cover $\mathfrak{R}_{p}(S)$, we apply a binary search strategy. For a candidate value $m$, we design a procedure that either produces a covering using $m+\lceil m/p\rceil]$, or certifies that no covering exists with $[m]$. This way, we identify the smallest value of $m$ for which $\mathfrak{R}_{p}(S)$ can be covered using the rounded covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, where $C=[m]$ and $D=[m-\beta]$.